OpenStudy (anonymous):

change the velocity of 40 km per hr to feet per second.

OpenStudy (anonymous):

\[40 \frac{ km }{ hr }\left( \frac{ 1 mile }{ 1.6 km } \right)\left( \frac{ 5280 ft. }{ 1 mile } \right)\left( \frac{ 1 hr }{ 60 \min } \right)\left( \frac{ 1 \min }{ 60 \sec. } \right)\]